# Large Language Models Assume People are More Rational than We Really are

### v3

@@ 3.1 Results
To evaluate whether LLMs are able to predict or simulate human risky choice decisions, we computed correlations and mean-squared errors (MSE) between LLMs’ responses and human decisions. We report both Pearson and Spearman correlation, but find that they are extremely similar and thus discuss them interchangeably. Additionally, we compared LLM responses to a classic model of rational choice: choosing the option with the highest expected value. In this case Pearson and Spearman correlation are identical because maximizing expected value results in a binary response. Due to space considerations we focus here on the first of our three tasks — predicting individual choices — because the results were extremely similar across the three tasks. Results from predicting the proportion of human choices and acting as a human appear in Appendix B.
@@
We found that LLMs with zero-shot prompting are not well-aligned with human decisions. Table 1 shows the model predictions of human behavior against actual human choices, with GPT-4-Turbo performing best with a correlation of 0.60 and an MSE of 0.13. These models are also not well-aligned with rational decision-making. Table 2 shows that the human correlation with the maximum expected value is 0.48, while even GPT-4-Turbo and GPT-4o only achieve correlations of 0.41 and 0.28, with MSEs of 0.27 and 0.36 respectively. [^1] In the zero-shot case, we ran a single completion for GPT-4o but multiple for GPT-4-Turbo, which likely resulted in GPT-4o being less correlated with human choices due to variance in sampling. In an effort to determine exactly what LLMs are doing, we fit a wider range of behavioral models to the responses of GPT-4-Turbo. This analysis suggests that the model does produce responses consistent with maximum expected value on a subset of choice problems, but also completely neglects probability information a significant proportion of the time (see Appendix D).
@@
We found all LLMs with CoT prompting assume people act rationally, which is not aligned with actual human decisions. As shown in Tables 1 and 2, even Llama3-8B achieves a correlation with maximum expected value of 0.57, while humans only achieve a correlation of 0.48. This correlation rises as model capabilities improve. Llama3-8B obtains a correlation of 0.57 with an MSE of 0.22; Llama3-70B obtains a correlation of 0.80 with an MSE of 0.1; Claude 3 Opus obtains a correlation of 0.76 with an MSE of 0.12, while GPT-4-Turbo (which best predicted people) and GPT-4o obtain correlations of 0.93 and 0.94 with MSEs of 0.03 and 0.02. The same patterns held in the tasks asking for aggregate behavior or for LLMs to act as humans, although the aggregate behavior task had slightly reduced correlations with expected value (see Appendix B for details).
@@
Although all the LLMs we investigated claim to be aligned with human preferences during the training, our empirical results suggest that aligning with “perfect” human data leads LLMs to assume humans act more rationally than they actually do, particularly when chain-of-thought prompting is used. In these settings, LLMs correlated more highly with maximizing expected value than with human choices, providing evidence of a gap between the implicit model of human decision-making assumed by LLMs and actual human behavior.
4 Inverse modeling: Inferring people’s preferences from their choices
While forward modeling predicts a person’s decision using established utilities, inverse modeling uses the same underlying decision-making models to infer a person’s utilities based on their made decisions, providing a complementary setting to evaluate the assumptions of LLMs. We adapt a task from a psychology experiment conducted by Jern et al. (2017).
@@ 4 Inverse modeling: Inferring people’s preferences from their choices
In the experiment, human participants were asked to place 47 observed decisions of others into a rank ordering indicating how much each decision suggested that the decision-maker prefers a target item. We used this paradigm rather than one that focuses on measuring the inferred preferences for each individual decision (e.g., Lucas et al., 2014), because such rankings provide more information about the implicit model assumed by the ranker. Through these rankings, Jern et al. (2017) found that humans ascribe almost perfectly rational decision-making models to the people they observe.
We found that GPT-4 Turbo could not provide a valid output ranking 47 choices at once. Instead, we obtained rankings by measuring pairwise comparisons between ($47\choose 2$ pairs of decisions. Pairwise outputs were limited to {stronger, weaker, tie}, and were aggregated across decisions to form a ranking. Ties were discouraged in the prompt to capture small differences between decisions.
@@
Each decision consisted of up to five unique items including the target item. Decisions were made between up to five distinct options. Items could be featured in multiple options within a decision, but never more than once in the same option. The target item was always in at least one option, and was always part of the chosen option. Observers did not know anything about the values of individual items — instead, relative utilities were inferred after observing the decision maker’s choice. Thus, in this paradigm, the remaining four items are equivalently exchangable; choosing {target item $X$} over {item $A$, item $B$} should yield same same inferred level of preference as choosing {target item $X$} over {item $C$, item $D$}. The 47 decisions were structurally unique, yielding coverage over all major decision types within this space. A full list of decisions is in Appendix G.
@@
Inverse decision-making models developed by psychologists to explain how people infer the preferences of others typically first specify a forward model and then infer preferences by applying Bayesian inference (Baker et al., 2009; Lucas et al., 2014; Jara-Ettinger et al., 2020; Ho and Griffiths, 2021). The forward model is normally a “noisy” version of a rational model, where options with greater utility are selected with higher probability. For example, given the utilities a decision-maker assigns to each item $\mathbf{u}$ and the attributes (items) within each of the $n$ options $\mathbf{A}=\{a_{1},\dots,a_{n}\}$, a standard model based on Luce (1959) assumes that the probability of choosing option $o_{j}$ in choice $c$ follows
@@
We constructed the prompts based on the original scripts and text instructions given to participants in the human experiment, adapted to pairwise comparisons instead of ranking 47 decisions at once. We also removed physical details of the experiment (e.g., the decisions were printed on cards with colors to represent items, while we describe items with natural language). The prompt first introduces the context (either candy or electric shocks), describes the pair of observed decisions, and concludes with the request to select the choice that more strongly suggests that the decision-maker prefers the target item. We also included additional experiment clarifications present in the original human experiment, as well as instructions for structuring the outputs (e.g., chain-of-thought) if applicable. To mitigate effects from LLMs’ positional bias (Wang et al., 2023) and any potential context biases related to item descriptions, we shuffled the individual contexts we assigned to each item (e.g., black vs. blue candy). We also shuffled the order that decisions appear in the pair, shuffled the order of options appearing within the decisions, and shuffled the order of items listed within each option. A prompt example is in Appendix F.
@@ 4.1 Results
We also observe that LLMs’ correlations with human inferences are consistently higher than their correlations with rational models. This is especially true in the negative context where humans are less consistent with the rational model (e.g., GPT-4o with CoT has a 0.87 correlation with humans, compared to a 0.74 correlation with the highest rational model). Thus, although LLMs typically assume rationality, when people’s inferences diverge from those of rational models, LLMs’ inferences are closer to humans. This could be explained by LLMs sharing some heuristic strategies with humans, but future investigations would be required to verify this. We provide scatterplots illustrating the patterns of responses in Appendix E.
Table 3: Spearman correlations between inverse decision rankings made by LLMs / humans and predictions from rational models. LLMs that most highly correlate with humans in each setting are in bold. Correlation coefficients with absolute value $\geq 0.3$ are statistically significant at $\alpha=0.05$, and $\geq 0.47$ at $\alpha=0.001$.
@@ 5 Discussion
We conducted an extensive evaluation of how LLMs assume people make decisions. In our forward modeling experiments, we found that LLMs struggle to predict or simulate human behavior in a simple risky choice setting, assuming that people make decisions more rationally than we actually do. We connect this to a previous finding in psychology — that people model others as more rational than they are — in order to explain why people think LLMs produce human-like behavior when making decisions. Then in our inverse modeling experiments, we find that LLMs also assume people act rationally when reasoning backwards from observed actions to internal utilities, aligning with how humans make inferences about others’ choices. Thus, LLMs seem to adopt a consistent model of human decision-making across forward and inverse modeling — one that assumes people act more rationally than we actually do.
In the remainder of this section, we cover broader impacts related to aligning LLMs and simulating humans using LLMs, discuss limitations to our work, and conclude with some notable future directions including the potential for cognitive science to help understand LLM alignment.
@@
The psychology literature shows that there is a dichotomy between how people make decisions and how we expect others to make decisions. How should alignment be defined when these are different? Existing frameworks that focus on safe and useful deployments (e.g., Bommasani et al., 2021; Askell et al., 2021) may prioritize aligning with our expectations, but there are also many merits to having models behave like us (e.g., Park et al., 2022; Shaikh et al., 2024). We believe a reasonable answer to this is to separate alignment into two sub-cases: alignment with human expectations and alignment with human behavior, and to train separate models that fulfill each objective. Models aligned with human expectation should shed human tendencies such as resource-rationality, i.e. sacrificing quality to reduce effort (Evans et al., 2015; Alanqary et al., 2021; Lieder and Griffiths, 2020), while models designed to simulate humans should retain them. By recognizing the difference between people’s expectations and behavior, we provide support for developing more specific alignment objectives grounded in insights from social science.
We hypothesize that certain training paradigms may be more suited towards aligning to human expectations, while others favor alignment with human behavior. For instance, high quality written responses used in Supervised Fine-Tuning may teach the LLM to mimic the original human writers, aligning outputs towards human behavior. On the other hand, when humans provide preferences for RLHF in a chat setting, their judgement might reflect what the human rater expects from the LLM.
@@
There is a growing literature investigating whether we can use LLMs to simulate humans for various applications (Park et al., 2023), such as acting as mock participants in human studies (Argyle et al., 2023; Aher et al., 2022; Hämäläinen et al., 2023), collecting public opinion (Chu et al., 2023; Kim and Lee, 2023; Sun et al., 2024), and helping provide realistic reactions to assist people’s communication (Liu et al., 2023; Shaikh et al., 2024; Lin et al., 2024; Shin and Kim, 2023). As our experiments show that LLMs make decisions more rationally than people do, and also predict people’s decisions to be more rational than they are, current LLMs may be fundamentally misaligned for the task of simulating human choices. Developing policy recommendations or even designing futher experiments based on the choices that LLMs make may be misleading — similar to the concerns about the use of an overly rational “homo economicus” originally raised by researchers in behavioral economics (Tversky and Kahneman, 1974; Kahneman and Tversky, 1979). Instead, we believe different training paradigms must be created to accomodate the divergent training goals of human expectation and behavior.
@@
Our experiments focused on a subset of all models and tasks for which we could evaluate models’ rationality. This subset was carefully selected to be representative of the existing literature, but could be expanded upon in future research. In particular, our experiments used controlled, abstract domains that cannot guarantee generalization to real-world contexts. While this is the same challenge faced by all human experiments, it is uniquely challenging for work done on LLMs due to their black-box nature and sensitivity to input prompts (Sclar et al., 2023). Furthermore, like all human data, the psychological datasets that we compare LLMs against were potentially subject to sampling bias and it is unknown whether they fully represent the true distribution of human rational choice or inference.
Another limitation is that we only use a simple simulation paradigm when testing LLMs’ capabilities in the risky choice experiments. While we rely on sampling temperature to create variance between individual participant samples, other works have used more advanced methods such as adding personal details, demographics, or traits for more realistic variation (Zhou et al., 2024b; Hu and Collier, 2024). Instead, our work evaluates the simplest LLM simulation method at scale, and the choices13k dataset did not contain such information regarding subjects, making a representative inclusion of these features impossible. Future work could conduct a more comprehensive analysis of different simulation methods and whether they alter LLMs’ implicit decision-making models.
@@
One peculiar observation that we found in the inverse modeling results is that LLMs’ inferences were consistently more correlated with people than they were with rational models. This suggests that LLMs may potentially capture aspects of human behavior that are not present in existing theoretical models. While we have focused on using theories and paradigms from psychology to analyze LLMs, there may also be opportunities to use LLMs to refine existing theories about people. More generally, our results show how studies and theories originating in psychology and cognitive science can help quantify the behavior of LLMs. These fields offer many other opportunities to compare the behavior of LLMs with humans, helping us towards understanding (both) these complex yet extremely capable systems.

### v4

@@ 3.1 Results
To evaluate whether LLMs are able to predict or simulate human risky choice decisions, we computed correlations and mean-squared errors (MSE) between LLMs’ responses and human decisions. Specifically, we compute the correlations across proportions where the LLM predicts that one gamble will be selected over the other. We report both Pearson and Spearman correlation, but they are extremely similar and we discuss them interchangeably. We also compared LLM responses to a classic model of rational choice: choosing the option with the highest expected value. [^1] Here, Pearson and Spearman correlation are identical as maximizing EV results in a binary response. We focus here on the first of our three tasks — predicting individual choices — because results were similar across the three tasks. Results from predicting the proportion of human choices and simulating decisions are in Appendix B.
@@
We found that LLMs with zero-shot prompting are not well-aligned with human decisions. Table 1 shows the model predictions of human behavior against real human choices, where GPT-4-Turbo performed best with a correlation of 0.60 and an MSE of 0.13. LLMs are also not well-aligned with rational decision-making. Table 2 shows that the human correlation with the maximum expected value is 0.48, while GPT-4-Turbo and 4o only achieve correlations of 0.41 and 0.28, with MSEs of 0.27 and 0.36 respectively. [^2] In the zero-shot case, we ran a single completion for GPT-4o but multiple for GPT-4-Turbo, which likely resulted in GPT-4o being less correlated with human choices due to variance in sampling.
LLMs using zero-shot prompts sometimes ignore probabilities completely.
To determine exactly what LLMs are doing in the zero-shot case, we fit 18 behavioral models to the responses of GPT-4-Turbo to examine its behavior. These included heuristic models (He et al., 2022), where people are thought to use mental shortcuts to make decisions, counterfactual models (He et al., 2022), which appeal to constructs like regret and disappointment, and subjective Expected Utility models, which assume that quantities involved (money and probability) are perceived or treated subjectively. This includes many of the most influential models including Prospect Theory (Kahneman and Tversky, 1979) and the model proposed by (Peterson et al., 2021), MOT.
In Appendix D, we provide the full analysis with these models. The best interpretation we found uses MOT, where the model fit a mixture of two probability weighting functions — one linear (matching humans), and one approximating a flat line. This suggests that the LLM often expected people to be rational, but completely ignores all probabilities (i.e., weighs them equally) in other cases.
@@
We found all LLMs with CoT prompting assume people act rationally, which is not aligned with actual human decisions. As shown in Tables 1 and 2, even Llama3-8B achieves a correlation with maximum expected value of 0.57, while humans only achieve a correlation of 0.48. This correlation rises as model capabilities improve: Llama3-8B obtains a correlation of 0.57 with an MSE of 0.22; Llama3-70B obtains a correlation of 0.80 with an MSE of 0.1; Claude 3 Opus obtains a correlation of 0.76 with an MSE of 0.12, while GPT-4-Turbo (which best predicted people) and GPT-4o obtain correlations of 0.93 and 0.94 with MSEs of 0.03 and 0.02. The same patterns held in the tasks asking for aggregate behavior or for LLMs to act as humans, although the aggregate behavior task had slightly reduced correlations with expected value (see Appendix B). In contrast, we also prompted GPT-4o to make the same predictions for monkeys — which were less rational, but actually closer to human choices (see Appendix E.3).
@@
Although all the LLMs we investigated claim to be aligned with human preferences during the training, our empirical results suggest that these LLMs assume humans act more rationally than they actually do, particularly when CoT prompting is used. In these settings, LLMs correlated more highly with maximizing expected value than with human choices, demonstrating a gap between the implicit model of human decision-making assumed by LLMs and actual human behavior.
4 Inverse modeling: Inferring people’s preferences from choices
While forward modeling allows us to predict someone’s decision given established utilities, inverse modeling is the process of inferring someone’s utilities via the decisions they make. Because forward and inverse modeling share the same theoretical framework, such inferences provide a complementary setting to evaluate the decision-making models that LLMs ascribe to humans. To formally measure these inferences, we adapt a psychology experiment from Jern et al. (2017).
@@ 4 Inverse modeling: Inferring people’s preferences from choices
In the experiment, participants observed 47 decisions made by another person (the observee). In each decision, the observee is shown $n$ groups of items $\{g_{1},\dots,g_{n}\}$ and choose the group $g_{p}$ which they prefer. Groups are populated with five distinct types of items, denoted $A,B,C,D,X$. For example, a particular decision could consist of two groups, $g_{1}=\{X\}$ and $g_{2}=\{A,B\}$, with $g_{p}=g_{1}$. Since the participant chose $g_{1}$, they prefer obtaining $X$ over obtaining both $A$ and $B$.
Participants were asked to rank the 47 observed decisions (which all contained item $X$) based on how much the decision suggested the observee has a high preference for $X$. For instance, the decision $(g_{1}=\{X\},g_{2}=\{A,B\},g_{p}=g_{1})$ shows a higher preference for $X$ than the naive decision $(g_{1}=\{X\},g_{p}=g_{1})$. Participants did not know the utilities of items. Thus, they were only able to determine the observee’s preferences based on features such as the number of items in each group.
The 47 decisions covered the most commonly distinguishable decision structures across items $A,B,C,D,X$, under the constraint that each item is included at most once per group. For simplicity, only decisions where $X$ is part of the preferred group $g_{p}$ were used. We provide a ranking of a subset of close decisions in Table 3, and a full list of 47 decisions in Appendix H.
Table 3: A subset of the 47 decisions that are closely ranked, ordered by average human ranking. $g_{p}=g_{1}$ for all rows. Utilities for items are positive.
shows less of a preference for X
$g_{1}=\{X\},$ $g_{2}=\{A\}$
$g_{1}=\{X,A,B\},$ $g_{2}=\{A,B,C\},$ $g_{3}=\{A,B,D\}$
$g_{1}=\{X,A\},$ $g_{2}=\{A,B\},$ $g_{3}=\{A,C\},$
$g_{1}=\{X,A\},$ $g_{2}=\{A,B\},$ $g_{3}=\{A,C\},$ $g_{4}=\{A,D\}$
$g_{1}=\{X\},$ $g_{2}=\{A\},$ $g_{3}=\{B\}$
shows more of a preference for X
We found that GPT-4 Turbo could not provide a valid output ranking 47 choices at once. Thus, for all LLMs, we obtained rankings by asking the LLM to perform pairwise comparisons across ($47\choose 2$ pairs of decisions. Pairwise outputs were limited to {stronger, weaker, tie}, and were aggregated across decisions to form a ranking. Ties were discouraged to capture small differences between decisions.
@@
Observers did not know anything about the values of individual items — instead, relative utilities were inferred after observing the decision maker’s choice. Thus, in this paradigm, the remaining four items are equivalently exchangable; choosing {target item $X$} over {item $A$, item $B$} should yield same same inferred level of preference as choosing {target item $X$} over {item $C$, item $D$}. The 47 decisions were structurally unique, yielding coverage over all major decision types within this space. A full list of decisions is in Appendix H.
@@
Through these rankings, Jern et al. (2017) found that humans ascribe almost perfectly rational decision-making models to the people they observe.
@@
Inverse decision-making models developed by psychologists to explain how people infer the preferences of others typically first specify a forward model and then infer preferences by applying Bayesian inference (Baker et al., 2009; Lucas et al., 2014; Jara-Ettinger et al., 2020; Ho and Griffiths, 2021). The forward model is normally a “noisy” version of a rational model, where options with greater utility are selected with higher probability. For example, given the utilities a decision-maker assigns to each item $\mathbf{u}$ and the items in each of the $n$ options $\mathbf{A}=\{a_{1},\dots,a_{n}\}$, a standard model based on Luce (1959) assumes the probability of choosing option $o_{j}$ in choice $c$ is
@@
We constructed prompts based on the original scripts and text instructions given to participants in the human experiment, adapted to pairwise comparisons instead of ranking 47 decisions at once. We also removed physical details of the experiment (e.g., the decisions were printed on cards with colors to represent items, while we describe items with natural language). The prompt first introduces the context (either candy or electric shocks), describes the pair of observed decisions, and concludes with the request to select the choice that more strongly suggests that the decision-maker prefers the target item. We also included additional clarifications present in the original human experiment, as well as instructions for structuring the outputs (e.g., chain-of-thought) if applicable. To mitigate effects from LLMs’ positional bias (Wang et al., 2023) and any potential context biases related to item descriptions, we shuffled the individual contexts we assigned to each item (e.g., black vs. blue candy). We also shuffled the order that decisions appear in the pair, shuffled the order of options within the decisions, and shuffled the order of items within each option. See Appendix G for examples.
@@ 4.1 Results
We also observe that LLMs’ correlations with human inferences are consistently higher than their correlations with rational models. This is especially true in the negative context where humans are less consistent with the rational model (e.g., GPT-4o with CoT has a 0.87 correlation with humans, compared to a 0.74 correlation with the highest rational model). Thus, although LLMs typically assume rationality, when people’s inferences diverge from those of rational models, LLMs’ inferences are closer to humans. This could be explained by LLMs sharing some heuristic strategies with humans, a topic for future investigation. Scatterplots showing patterns of responses are in Appendix F.
[…]
@@ 5 Discussion
We conducted an extensive evaluation of how LLMs assume people make decisions. In our forward modeling experiments, we found that LLMs struggle to predict or simulate human behavior in a classic risky choice setting, assuming that people make decisions more rationally than we actually do. We connect this to a previous finding in psychology — that people model others as more rational than they are — in order to explain why people think LLMs produce human-like behavior when making decisions. Then in our inverse modeling experiments, we find that LLMs also assume people act rationally when reasoning backwards from observed actions to internal utilities, aligning with how humans make inferences about others’ choices. Thus, LLMs seem to adopt a consistent model of human decision-making across forward and inverse modeling — one that assumes people act more rationally than we actually do.
@@
The psychology literature shows that there is a dichotomy between how people make decisions and how we expect others to make decisions. How should alignment be defined when these are different? Existing frameworks that focus on safe and useful deployments (e.g., Bommasani et al., 2021; Askell et al., 2021) may prioritize aligning with our expectations, but there are also many merits to having models behave like us (e.g., Park et al., 2022; Shaikh et al., 2024). We believe a reasonable answer to this is to separate alignment into two sub-cases: alignment with human expectations and alignment with human behavior, and to train separate models that fulfill each objective. Models aligned with human expectation should shed human tendencies such as resource-rationality, i.e. sacrificing quality to reduce effort (Evans et al., 2015; Alanqary et al., 2021; Lieder and Griffiths, 2020), while models designed to simulate humans should retain them. By recognizing the difference between people’s expectations and behavior, we provide support for developing more specific alignment objectives grounded in social science.
We hypothesize that certain training paradigms may be more suited towards aligning to human expectations, while others favor alignment with human behavior. For instance, high-quality written responses used in Supervised Fine-Tuning may teach the LLM to mimic the original human writers, aligning outputs towards human behavior. On the other hand, when humans provide preferences for RLHF in a chat setting, their judgement might reflect what the human rater expects from the LLM.
@@
There is a growing literature investigating whether we can use LLMs to simulate humans for various applications (Park et al., 2023), such as acting as mock participants in human studies (Argyle et al., 2023; Aher et al., 2022; Hämäläinen et al., 2023), collecting public opinion (Chu et al., 2023; Kim and Lee, 2023; Sun et al., 2024), and helping provide realistic reactions to assist people’s communication (Liu et al., 2023; Shaikh et al., 2024; Lin et al., 2024; Shin and Kim, 2023). As our experiments show that LLMs make decisions more rationally than people do, and also predict people’s decisions to be more rational than they are, current LLMs are fundamentally misaligned for the task of simulating human choices. Developing policy recommendations or designing experiments based on LLM-simulated choices may be misleading — similar to concerns about the use of an overly rational “homo economicus” originally raised by researchers in behavioral economics (Tversky and Kahneman, 1974; Kahneman and Tversky, 1979).
@@
Our experiments focused on a subset of all tasks for which we could evaluate models’ rationality. This subset was carefully selected to be representative of the classic literature, but could be expanded upon in future research. In particular, our experiments used controlled, abstract domains that do not guarantee generalization to real-world contexts. While this is the same challenge faced by all human experiments, it is uniquely challenging for work done on LLMs due to their black-box nature and sensitivity to input prompts (Sclar et al., 2023). Furthermore, like all human data, the psychological datasets that we compare LLMs against were potentially subject to sampling bias and it is unknown whether they fully represent the true distribution of human choices or inferences.
Another limitation is that we use simple simulation paradigms for testing LLMs’ capabilities. While we conduct an ablation using demographic prompts, other methods such as adding personal details or traits for more realistic variation have been proposed (Zhou et al., 2024b; Hu and Collier, 2024). The choices13k dataset does not contain such information about subjects, making a representative inclusion of these features impossible. However, future work could conduct a more comprehensive analysis of different simulation methods to see if they alter LLMs’ implicit decision-making models.
@@
A peculiar observation from the inverse modeling setting is that LLMs’ inferences were consistently more correlated with people than they were with rational models. This suggests that LLMs may potentially capture aspects of human behavior that are not present in existing theoretical models. While we have focused on using theories and paradigms from psychology to analyze LLMs, there may also be opportunities to use LLMs to refine existing theories about people. More generally, our results show how studies and theories originating in psychology and cognitive science can help quantify the behavior of LLMs. These fields offer many more opportunities to compare the behavior of LLMs with humans, helping us towards understanding (both) these complex yet extremely capable systems.